# Isolated Causal Effects of Natural Language

### v1

@@ 1.1.5 Confidence intervals for $\widehat{\tau}_{DR}$
Asymptotic normality is established using the CLT (Kennedy,, 2023):
@@
1.2 OVB metrics
Following Chernozhukov et al., (2024), we can define a “short” version of our isolated effect estimand as the difference between $g(1,\cdot)-g(0,\cdot)$ where we use the “short” representation of the non-focal language, $a^{\mathsf{c}}_{s}(X)$, in place of the true representation $a^{\mathsf{c}}(X^{\ast})$:
@@ 1.2 OVB metrics
Using the proof in Appendix 1.1.1, we also have that
@@
This allows us to align our estimand with Chernozhukov et al., (2024), where $g_{s}$ is the short outcome model ($g(a^{\prime},a_{s}^{\mathsf{c}}(X))$ in our setting) and $\alpha_{s}$ are the short Riesz representer weights ($\gamma(a^{\prime},a_{s}^{\mathsf{c}}(X))$ in our setting). Then it follows directly that:
@@
1.3 OVB estimators
Following the procedure in Chernozhukov et al., (2024), we construct debiased estimators for $\sigma^{2}$ and $\nu^{2}$.
@@ 1.3 OVB estimators
2 ADDITIONAL RESULTS
(a) Isolated effect of female.
(b) Isolated effect of filler.
(c) Isolated effect of sexual.
In this section, we provide and discuss isolated effect estimates for additional interventions on the Amazon dataset (Figure 5). As is the case for the results in the main paper, we see here that as the dimensionality of $a_{s}^{\mathsf{c}}(X)$ increases, fidelity improves (i.e., the fidelity metric decreases) while overlap becomes worse (i.e., the overlap metric increases). [^3] The last couple of dimensions for female are an exception to this, where overlap slightly improves—this may be by chance due to better regularization in the classifier models. The robustness value suggests overall improvement with increasing dimensionality, though we again note that this may not be the case for some datasets where overlap violations outweigh fidelity improvements.
@@ 2 ADDITIONAL RESULTS
3 EXPERIMENTS
3.1 Data
@@
Table 1: Composition of data splits and licensing information.
3.2 Language representation implementation
Language Library Version
LIWC Python liwc 0.5.0
Empath Python empath 0.89
MPNet embedding Python HuggingFace sentence-transformers (all-mpnet-base-v2) 2.2.2
SenteCon Python sentecon 0.1.9
GPT-3.5 prompting - - -
@@ 3.2 Language representation implementation
To implement our lexicons, we use the third-party liwc Python library and the empath library released by its creators. MPNet embeddings are obtained via the HuggingFace sentence-transformers library using the pre-trained model all-mpnet-base-v2. SenteCon-LIWC and SenteCon-Empath representations are obtained using the sentecon library released by its creators. Finally, LLM (GPT-3.5) prompting covariates are taken directly from the SvT dataset released by Dhawan et al., (2024). Additional technical details are provided in Table 2.
3.3 Model details and hyperparameters
All outcome models and $a(X)$ classifiers are implemented using the scikit-learn Python library (version 1.3.0). Gradient boosting models use a subsample proportion of 0.7, i.e., 70% of training samples are used to fit the individual base learners.
@@ 3.3 Model details and hyperparameters
3.4 OVB lower bound analysis (computing $C_{Y}$ and $C_{D}$)
@@
Then following Chernozhukov et al., (2024), the explanatory power lost from this omitted information can be computed explicitly as $C_{Y}$ and $C_{D}$:

### v3

@@ A.1.5 Confidence Intervals for $\widehat{\tau}_{DR}$
Asymptotic normality is established using the CLT (Kennedy, 2024):
@@
A.2 OVB Metrics
Following Chernozhukov et al. (2024), we can define a “short” version of our isolated effect estimand as the difference between $g(1,\cdot)-g(0,\cdot)$ where we use the “short” representation of the non-focal language, $a^{\mathsf{c}}_{s}(X)$, in place of the true representation $a^{\mathsf{c}}(X^{\ast})$:
@@ A.2 OVB Metrics
Using the proof in Appendix A.1.1, we also have that
@@
This allows us to align our estimand with Chernozhukov et al. (2024), where $g_{s}$ is the short outcome model ($g(a^{\prime},a_{s}^{\mathsf{c}}(X))$ in our setting) and $\alpha_{s}$ are the short Riesz representer weights ($\gamma(a^{\prime},a_{s}^{\mathsf{c}}(X))$ in our setting). Then it follows directly that:
@@
A.3 OVB Estimators
Following the procedure in Chernozhukov et al. (2024), we construct debiased estimators for $\sigma^{2}$ and $\nu^{2}$.
@@ A.3 OVB Estimators
Appendix B Additional Results
B.1 Additional Interventions (Linear Amazon Outcome)
Figure 5: Isolated causal effects of linguistic attributes on helpfulness in the Amazon dataset (linear semi-synthetic outcome). Error bars correspond to 95% confidence intervals.
In this section, we provide and discuss isolated effect estimates for additional interventions on the Amazon dataset (Figure 5). As is the case for the results in the main paper, we see here that as the dimensionality of $a_{s}^{\mathsf{c}}(X)$ increases, fidelity improves (i.e., the fidelity metric decreases) while overlap becomes worse (i.e., the overlap metric increases). [^4] The last several dimensions for female are an exception to this, where overlap slightly improves—this may be by chance due to better regularization in the classifier models. The robustness value suggests overall improvement with increasing dimensionality, though we again note that this may not be the case for some datasets where overlap violations outweigh fidelity improvements.
@@ B.1 Additional Interventions (Linear Amazon Outcome)
B.2 Nonlinear Amazon Outcome
Figure 6: Isolated causal effects of linguistic attributes on helpfulness in the Amazon dataset (nonlinear semi-synthetic outcome). Error bars correspond to 95% confidence intervals.
In this section, we discuss results of isolated effect estimation on a more complex version of the Amazon dataset where the semi-synthetic outcome is a nonlinear function of $a(X),a^{\mathsf{c}}(X)$. The outcome in this setting differs from the one described in Section 4.1 only in that a nonlinear gradient boosting model is used to predict helpful vote count from $a(X),a^{\mathsf{c}}_{s}(X)$ instead of a linear regression model; this prediction is then noised to obtain the semi-synthetic outcome.
Using this new outcome, we follow the protocol described in Section 5.1 to obtain effect estimates for the two attributes featured in Figure 2: home and netspeak. During estimation, we use simple feedforward neural networks with no more than 3 layers to fit our importance weight and outcome models.
The results of these additional experiments are consistent with those included in the main paper. For both interventions, we observe that the point estimate of the isolated effect generally grows closer to the ground truth as the number of dimensions increases (i.e., as the number of omitted variables decreases). Likewise, the behavior of the fidelity and overlap metrics $\widehat{\sigma}^{2}$ and $\widehat{\nu}^{2}$ remains consistent with expectations: as dimensionality increases, so does $\widehat{\nu}^{2}$, while $\widehat{\sigma}^{2}$ decreases. Occasionally slight variability in these trends appears, as we would expect from the noisiness of estimation with more complex models.
Turning to robustness, we see that for home, robustness value generally increases with the number of dimensions, suggesting that gains in model fidelity outweigh losses in overlap. For netspeak, robustness value remains about the same up until 6 features, then decreases sharply. This coincides with the effect point estimate starting to move away from the ground truth, though the estimate’s confidence intervals do still contain the true effect. This suggests that the worsening overlap outweighs gains in model fidelity and that for this effect, the optimal non-focal language representation $a_{s}^{\mathsf{c}}(X)$ may contain only 6 features.
Finally, we note that once the final feature is added for netspeak, $\widehat{\nu}^{2}$ sharply increases, signaling much worse overlap. This is an interesting illustration of how soft overlap violations can occur once sufficient information is contained in $a_{s}^{\mathsf{c}}(X)$ such that the model can fully predict a(X).
Appendix C Experiments
C.1 Data
Table 1: Composition of data splits and licensing information.
@@ C.1 Data
C.2 Language Representation Implementation
@@
Language Library Version Model
LIWC Python liwc 0.5.0 -
Empath Python empath 0.89 -
SenteCon Python sentecon 0.1.9 -
BERT embedding Python transformers 4.32.1 bert-base-uncased
RoBERTa embedding Python transformers 4.32.1 roberta-base
MiniLM embedding Python sentence-transformers 2.2.2 all-MiniLM-L6-v2
MPNet embedding Python sentence-transformers 2.2.2 all-mpnet-base-v2
GPT-3.5 prompting - - - gpt-3.5-turbo
To implement our lexicons, we use the third-party liwc Python library and the empath library released by its creators. SenteCon-LIWC and SenteCon-Empath representations are obtained using the sentecon library released by its creators. BERT and RoBERTa embeddings are obtained via the HuggingFace transformers library using the pre-trained models bert-base-uncased and roberta-base, respectively. MPNet and MiniLM embeddings are obtained via the HuggingFace sentence-transformers library using the pre-trained models all-mpnet-base-v2 and all-MiniLM-L6-v2, respectively. Finally, LLM (GPT-3.5) prompting covariates are taken directly from the SvT dataset released by Dhawan et al. (2024). Additional technical details are provided in Table 2.
C.3 Model Details and Hyperparameters
All outcome models and $a(X)$ classifiers are implemented using the scikit-learn Python library (version 1.3.0). Gradient boosting models use a subsample proportion of 0.7, i.e., 70% of training samples are used to fit the individual base learners. Neural networks used for outcome models in the nonlinear Amazon setting are implemented with the MLPRegressor class and tuned over the following possible layer counts and sizes: (128,), (128, 128), (128, 256, 128).
@@ C.3 Model Details and Hyperparameters
C.4 OVB Lower Bound Analysis (Computing $C_{Y}$ and $C_{D}$)
@@
Then following Chernozhukov et al. (2024), the explanatory power lost from this omitted information can be computed explicitly as $C_{Y}$ and $C_{D}$: